# Construction of graph-directed invariant sets of weak contractions on semi-metric spaces

### v1

@@ 2.5. Product space
Hereafter, we need to deal with product spaces $X^{q}$ and $\mathcal{F}(X)^{q}$. Elements in those product spaces are written in bold face.
@@
(i) If $(X,d)$ be a semi-metric space, then, $(X^{q},d_{\infty})$ is a semi-metric space.
@@
Hence, $d_{HP}^{\infty}({\bf A},{\bf C})\leq\Phi_{d_{HP}}\left(d_{HP}^{\infty}({\bf A},{\bf B}),d_{HP}^{\infty}({\bf B},{\bf C})\right)$. Hence, $\Phi_{d_{HP}^{\infty}}(u,v)\leq\Phi_{d_{HP}}(u,v)$. By Lemma 2.8, $(\mathcal{F}(X),d_{HP})$ satisfies the normal property and the regular property. Thus we have the normal property and the regular property for $(\mathcal{F}(X)^{q},d_{HP}^{\infty})$. ∎
@@ 3. Proof
Let ${\bf H}_{0}$ as in the proof of Lemma 3.1. Let ${\bf H}_{n}:=T^{n}({\bf H}_{0})$ for $n\geq 1$. Then, we see that ${\bf H}_{n}\in\mathcal{F}(X)^{q}$ by induction in $n$. By Lemma 2.12, $(\mathcal{F}(X)^{q},d_{HP}^{\infty})$ is a normal and regular symmetric pre-metric space. By Lemma 2.14, $T:\mathcal{F}(X)^{q}\to\mathcal{F}(X)^{q}$ is a $\phi$-contraction on $(\mathcal{F}(X)^{q},d_{HP}^{\infty})$. By Lemma 2.5, $({\bf H}_{n})_{n}$ is a Cauchy sequence in $(\mathcal{F}(X)^{q},d_{HP}^{\infty})$ and hence it is bounded. Hence, there exists $r_{0}>0$ such that $d_{HP}^{\infty}\left({\bf H}_{n},{\bf H}_{0}\right)<r_{0}$ for every $n\geq 1$.
It holds that $T$ is an isotone map, that is, $T({\bf A})\subset T({\bf B})$ holds if ${\bf A}\subset{\bf B}$. Therefore, ${\bf H}_{n+1}=T({\bf H}_{n})\supset{\bf H}_{n}$ for $n\geq 0$. Let ${\bf H}_{n}=:(H_{n,1},\dots,H_{n,q})$ for $n\geq 0$. Then, for each $i\in V$, $(H_{n,i})_{n}$ is an increasing sequence of subsets of $X$ and furthermore, $H_{n,i}\subset H_{0,i}(r_{0})$ for every $n\geq 0$. Let $H_{i}:=\cup_{n\geq 0}H_{n,i}$. Then, $H_{i}\subset H_{0,i}(r_{0})$ and $H_{i}\neq\emptyset$. Since $H_{0i}$ is bounded, by Lemma 2.2 (ii), $H_{0,i}(r_{0})$ is also bounded. Hence, $H_{i}$ is a bounded subset of $X$. Hence, $H_{i}\in\mathcal{F}(X)$. Let ${\bf H}:=(H_{1},\dots,H_{q})$. Let $T({\bf H})=:\left(T({\bf H})_{1},\dots,T({\bf H})_{q}\right)$. Then, by (1.1),
@@
In the proof of [4, Theorem 1], it is stated that the union is bounded. This is correct, however, there was no proof in [4]. The above proof is one way to fix it, and the strategy is similar to the proof of [3, Theorem 1].
For the Kuratowski non-compactness measure, let
@@
We first show that $T\left({\bf\overline{H}}\right)\subset{\bf\overline{H}}$. Let $y=(y_{1},\dots,y_{q})\in T\left({\bf\overline{H}}\right)$. Then, for each $i\in V$, there exists $j\in V$ and $e\in E_{i,j}$ such that $y_{i}\in f_{e}(\overline{H_{j}})$. Then, there exists a sequence $(x_{k})_{k}$ in $H_{j}$ such that $d(y_{i},f_{e}(x_{k}))\to 0,k\to\infty$. Since $T({\bf H})={\bf H}$, $f_{e}(H_{j})\subset H_{i}$. Hence, $y_{i}\in\overline{H_{i}}$. Hence, $y\in{\bf\overline{H}}$.
@@
$T\left({\bf\overline{H}}\right)_{i}=\bigcup_{j=1}^{q}\bigcup_{e\in E_{i,j}}f_{e}(\overline{H_{j}}),$
Since $\overline{H_{j}}$ is compact and $f_{e}$ is continuous, $f_{e}(\overline{H_{j}})$ is also compact. Since the union in the right hand side is a finite union, $T\left({\bf\overline{H}}\right)_{i}$ is also compact in $X$. By [1, Lemma 1], $X$ is Hausdorff. Hence, $T\left({\bf\overline{H}}\right)_{i}$ is closed in $X$. Since $T({\bf H})={\bf H}$, $H_{i}\subset T\left({\bf\overline{H}}\right)_{i}$. Thus we see that $\overline{H_{i}}\subset T\left({\bf\overline{H}}\right)_{i}$. ∎
@@ 4. Transform for semi-metric
The following functions are semi-metric transformer.
@@
(iii) $\Psi(t):=\left(\frac{t}{1+t}\right)^{\alpha}$, $\alpha>0$. The semi-metric in [9, Example 2.2] is obtained if $\alpha=4$ and $d$ is the Euclid distance.
@@
We denote the basic triangle function for $\widetilde{d}$ by $\Phi_{\widetilde{d}}$. Let $u,v\in[0,\Psi(\infty))$. Assume that $\widetilde{d}(x,z)\leq u$ and $\widetilde{d}(y,z)\leq v$. Since $\Psi$ is increasing, $d(x,z)\leq\Psi^{(-1)}(u)<\infty$ and $d(y,z)\leq\Psi^{(-1)}(v)<\infty$. Hence, $d(x,y)\leq\Phi_{d}\left(\Psi^{(-1)}(u),\Psi^{(-1)}(v)\right)$. Hence,
@@
(i) The assumption that $\Psi$ is continuous at 00 is not used for the statement of the normal property. The normal property concerns the global property of semi-metric.
(ii) It is easy to see that $\Psi\circ d$ is uniformly equivalent with $d$, so we can apply a general result in [7, Theorem 3.2] to the regular property. The regular property concerns the local property of semi-metric.
@@
Let $X=[0,1]$ and $d(x,y):=|x-y|^{\alpha}$, where $|x-y|$ is the Euclid distance and $\alpha>0$. Then, $d$ is a regular, normal and complete semi-metric on $X$. We consider a graph-directed set on $X$. Let $q=2$. Assume that $|E_{i,j}|=1$ for every $i,j\in\{1,2\}$. We denote $f_{e}$ for $e\in E_{i,j}$ by $f_{i,j}$. Let
@@
Let $A_{0}:=\{0\}$ and $B_{0}:=\{1\}$. Let ${\bf H}_{0}:=(A_{0},B_{0})$. Since $f_{1,1}(0)=0$ and $f_{2,2}(1)=1$, ${\bf H}_{0}\subset T({\bf H}_{0})$. Let
@@
Let $A_{n+1}:=g_{1,1}(A_{n})\cup g_{1,2}(A_{n})$ and $B_{n+1}=g_{2,1}(B_{n})\cup g_{2,2}(B_{n})$. Let $r(x):=1-x$.
We show that $r(A_{n})=B_{n}$ by induction in $n$. The case that $n=0$ is obvious. Assume that $r(A_{n})=B_{n}$. Then, since $r(r(x))=x$, $r\circ g_{1,1}\circ r=g_{2,2}$, $r\circ g_{1,2}\circ r=g_{2,1}$, it holds that $A_{n}=r(B_{n})$ and
@@
We show that $T^{n}({\bf H}_{0})=(A_{n},B_{n})$ by induction in $n$. The case that $n=0$ is obvious. Assume that $T^{n}({\bf H}_{0})=(A_{n},B_{n})$. Then, by (4.3),
@@
Let $K_{i},i=1,2$, be non-empty compact subsets of $X=[0,1]$ such that $K_{i}=g_{i,1}(K_{i})\cup g_{i,2}(K_{i}),i=1,2$. Let ${\bf K}:=(K_{1},K_{2})$. Then, $T({\bf K})={\bf K}$ and ${\bf K}=\overline{\cup_{n}T^{n}({\bf H}_{0})}$. We see that $r(K_{1})=K_{2}$.
Let $S(H):=\cup_{i,j\in\{1,2\}}f_{i,j}(H)$. Let $K$ be a non-empty compact subset of $[0,1]$ such that $K=S(K)$. Then, $K$ is a generalized Cantor set, and in terms of the base-7 expansion, we obtain that
@@
in particular, $f_{i,j}([0,1])$ are disjoint. Since each $f_{i,j}$ is injective, $f_{i,j}(K_{j})\subsetneq f_{i,j}(K_{1}\cup K_{2})\subset f_{i,j}([0,1])$. Therefore,
@@
We can give many examples. [20, Example 3] gives an example different from the above.
@@ 5. Graph-directed de Rham functional equation
In this section, we generalize the framework of the de Rham functional equation to the graph-directed setting. Functional equation indexed by a directed network was recently considered by [21]. Here we deal with the functional equation on $[0,1]$ indexed by a graph-directed finite graph in terms of the graph-directed invariant set and conjugate equations [6, 22]. For simplicity we focus on the case corresponding to $q=2$.
@@
We state several properties of compatible family $h$. Let
@@
Since $\left|I^{h}_{i}(i_{1},\dots,i_{n})\right|\leq\phi^{n}(1)$ and $\phi$ ia a comparison function,
@@
Let $\{f_{i,j}\}_{i,j}$ and $\{g_{i,j}\}_{i,j}$ be two compatible families on $[0,1]$ with common comparison function $\phi$ under the Euclid distance [^2] We can replace the Euclid metric with a different metric by the semi-metric transformer in the above section, as long as $\{f_{i,j}\}_{i,j}$ and $\{g_{i,j}\}_{i,j}$ are contractions in the sense of Matkowski under the new metric.. For each $i=1,2$, we define functions $\varphi_{i}$, on $D^{f}_{i}$ such that for $i_{1},\dots,i_{n}\in\{1,2\},x_{n}\in\{0,1\}$,
@@
In the same manner, we see that $(i_{k},j_{k})=(2,1)$ for $k\in\{\ell+1,\dots,n\}$ by induction in $k$, and $(x_{n},x_{n}^{\prime})=(1,0)$. By the compatibility condition (5.1) for $\{g_{i,j}\}$, it holds that
@@
On $[0,1]$, we put the topology induced by the Euclid metric.
@@
(iii) A pair of maps $(\varphi_{1},\varphi_{2})$ satisfies the following systems of functional equations:
@@
Assume that $x\in D^{h}_{i}\cap(0,1)$. Then, there exist $n\geq 1$ such that $x\in D^{h}_{i,n}\setminus D^{h}_{i,n-1}$, where we let $D^{h}_{i,0}:=\emptyset$. Then, there exist $i_{1},\dots,i_{n-1}\in\{1,2\}$ such that
@@
Second, we state connection of the functional equation with the graph-directed set. Put the product topology on $[0,1]^{2}$. Let
@@
Let ${\bf D}_{n}:=(D_{1,n},D_{2,n})$. By induction in $n$, we see that $T^{n}\left({\bf H}_{0}\right)={\bf D}_{n-1}$ for $n\geq 2$. By the arguments in Section 3,
@@
The discrete approximation is a natural approach for analysis for (5.6) and the framework of Theorem 1.1 aligns with the functional equation well.

### v2

@@ 2.5. Product space
Hereafter, we need to deal with product spaces $X^{q}$ and $\mathcal{F}(X)^{q}$. Elements in these product spaces are written in bold face.
@@
(i) If $(X,d)$ is a semi-metric space, then, $(X^{q},d_{\infty})$ is a semi-metric space.
@@
Hence, $d_{HP}^{\infty}({\bf A},{\bf C})\leq\Phi_{d_{HP}}\left(d_{HP}^{\infty}({\bf A},{\bf B}),d_{HP}^{\infty}({\bf B},{\bf C})\right)$. So, $\Phi_{d_{HP}^{\infty}}(u,v)\leq\Phi_{d_{HP}}(u,v)$. By Lemma 2.8, $(\mathcal{F}(X),d_{HP})$ satisfies the normal property and the regular property. Thus we have the normal property and the regular property for $(\mathcal{F}(X)^{q},d_{HP}^{\infty})$. ∎
@@ 3. Proof
Let ${\bf H}_{0}$ be as in the proof of Lemma 3.1. Let ${\bf H}_{n}:=T^{n}({\bf H}_{0})$ for $n\geq 1$. Then, we see that ${\bf H}_{n}\in\mathcal{F}(X)^{q}$ by induction on $n$. By Lemma 2.12, $(\mathcal{F}(X)^{q},d_{HP}^{\infty})$ is a normal and regular symmetric pre-metric space. By Lemma 2.14, $T:\mathcal{F}(X)^{q}\to\mathcal{F}(X)^{q}$ is a $\phi$-contraction on $(\mathcal{F}(X)^{q},d_{HP}^{\infty})$. By Lemma 2.5, $({\bf H}_{n})_{n}$ is a Cauchy sequence in $(\mathcal{F}(X)^{q},d_{HP}^{\infty})$ and hence it is bounded. Hence, there exists $r_{0}>0$ such that $d_{HP}^{\infty}\left({\bf H}_{n},{\bf H}_{0}\right)<r_{0}$ for every $n\geq 1$.
It holds that $T$ is an isotone map, that is, $T({\bf A})\subset T({\bf B})$ holds if ${\bf A}\subset{\bf B}$. Therefore, ${\bf H}_{n+1}=T({\bf H}_{n})\supset{\bf H}_{n}$ for $n\geq 0$. Let ${\bf H}_{n}=:(H_{n,1},\dots,H_{n,q})$ for $n\geq 0$. Then, for each $i\in V$, $(H_{n,i})_{n}$ is an increasing sequence of subsets of $X$ and furthermore, $H_{n,i}\subset H_{0,i}(r_{0})$ for every $n\geq 0$. Let $H_{i}:=\cup_{n\geq 0}H_{n,i}$. Then, $H_{i}\subset H_{0,i}(r_{0})$ and $H_{i}\neq\emptyset$. Since $H_{0i}$ is bounded, by Lemma 2.2 (ii), $H_{0,i}(r_{0})$ is also bounded. Hence, $H_{i}$ is a bounded subset of $X$. Thus, $H_{i}\in\mathcal{F}(X)$. Let ${\bf H}:=(H_{1},\dots,H_{q})$ and $T({\bf H})=:\left(T({\bf H})_{1},\dots,T({\bf H})_{q}\right)$. Then, by (1.1),
@@
In the proof in [5, Theorem 1], it is stated that the union is bounded. This is correct, however, there was no proof in [5]. The above proof is one way to fix it, and the strategy is similar to the proof in [4, Theorem 1].
For the Kuratowski measure of non-compactness, let
@@
We first show that $T\left({\bf\overline{H}}\right)\subset{\bf\overline{H}}$. Let $y=(y_{1},\dots,y_{q})\in T\left({\bf\overline{H}}\right)$. Then, for each $i\in V$, there exist $j\in V$ and $e\in E_{i,j}$ such that $y_{i}\in f_{e}(\overline{H_{j}})$. Then, there exists a sequence $(x_{k})_{k}$ in $H_{j}$ such that $d(y_{i},f_{e}(x_{k}))\to 0,k\to\infty$. Since $T({\bf H})={\bf H}$, $f_{e}(H_{j})\subset H_{i}$. Hence, $y_{i}\in\overline{H_{i}}$, and then, $y\in{\bf\overline{H}}$.
@@
$T\left({\bf\overline{H}}\right)_{i}=\bigcup_{j=1}^{q}\bigcup_{e\in E_{i,j}}f_{e}(\overline{H_{j}}).$
Since $\overline{H_{j}}$ is compact and $f_{e}$ is continuous, $f_{e}(\overline{H_{j}})$ is also compact. Since the union on the right hand side is a finite union, $T\left({\bf\overline{H}}\right)_{i}$ is also compact in $X$. By [2, Lemma 1], $X$ is Hausdorff. Hence, $T\left({\bf\overline{H}}\right)_{i}$ is closed in $X$. Since $T({\bf H})={\bf H}$, $H_{i}\subset T\left({\bf\overline{H}}\right)_{i}$. Thus we see that $\overline{H_{i}}\subset T\left({\bf\overline{H}}\right)_{i}$. ∎
@@ 4. Transform for semi-metric
The following functions are semi-metric transformers.
@@
(iii) $\Psi(t):=\left(\frac{t}{1+t}\right)^{\alpha}$, $\alpha>0$. The semi-metric in [10, Example 2.2] is obtained if $\alpha=4$ and $d$ is the Euclidean distance.
@@
We denote the basic triangle function for $\widetilde{d}$ by $\Phi_{\widetilde{d}}$. Let $u,v\in[0,\Psi(\infty))$. Assume that $\widetilde{d}(x,z)\leq u$ and $\widetilde{d}(y,z)\leq v$. Since $\Psi$ is increasing, $d(x,z)\leq\Psi^{(-1)}(u)<\infty$ and $d(y,z)\leq\Psi^{(-1)}(v)<\infty$. Hence, $d(x,y)\leq\Phi_{d}\left(\Psi^{(-1)}(u),\Psi^{(-1)}(v)\right)$. Thus,
@@
(i) The assumption that $\Psi$ is continuous at 00 is not used in the statement of normal property. The normal property concerns the global property of semi-metric.
(ii) It is easy to see that $\Psi\circ d$ is uniformly equivalent to $d$, so we can apply a general result in [8, Theorem 3.2] to the regular property. The regular property concerns the local property of semi-metric.
@@
Let $X=[0,1]$ and $d(x,y):=|x-y|^{\alpha}$, where $|x-y|$ is the Euclidean distance and $\alpha>0$. Then, $d$ is a regular, normal and complete semi-metric on $X$. We consider a graph-directed set on $X$. Let $q=2$. Assume that $|E_{i,j}|=1$ for every $i,j\in\{1,2\}$. We denote $f_{e}$ for $e\in E_{i,j}$ by $f_{i,j}$. Let
@@
We define finite sets $A_{n}$ and $B_{n}$, $n\geq 1$, inductively. Let $A_{0}:=\{0\}$ and $B_{0}:=\{1\}$. Let ${\bf H}_{0}:=(A_{0},B_{0})$. Since $f_{1,1}(0)=0$ and $f_{2,2}(1)=1$, ${\bf H}_{0}\subset T({\bf H}_{0})$. Let
@@
Let $A_{n+1}:=g_{1,1}(A_{n})\cup g_{1,2}(A_{n})$ and $B_{n+1}=g_{2,1}(B_{n})\cup g_{2,2}(B_{n})$ for $n\geq 0$. Let $r(x):=1-x$.
We show that $r(A_{n})=B_{n}$ by induction on $n$. The case that $n=0$ is obvious. Assume that $r(A_{n})=B_{n}$. Then, since $r(r(x))=x$, $r\circ g_{1,1}\circ r=g_{2,2}$, $r\circ g_{1,2}\circ r=g_{2,1}$, it holds that $A_{n}=r(B_{n})$ and
@@
We show that $T^{n}({\bf H}_{0})=(A_{n},B_{n})$ by induction on $n$. The case that $n=0$ is obvious. Assume that $T^{n}({\bf H}_{0})=(A_{n},B_{n})$. Then, by (4.3),
@@
Let $K_{i},i=1,2$, be non-empty compact subsets of $X=[0,1]$ such that $K_{i}=g_{i,1}(K_{i})\cup g_{i,2}(K_{i}),i=1,2$. Let ${\bf K}:=(K_{1},K_{2})$. Then,
$T({\bf K})={\bf K}=(f_{1,1}(K_{1})\cup f_{1,2}(K_{2}),f_{2,1}(K_{1})\cup f_{2,2}(K_{2}))=\overline{\cup_{n}T^{n}({\bf H}_{0})}.$ (4.4)
We see that $r(K_{1})=K_{2}$.
Let $S(H):=\cup_{i,j\in\{1,2\}}f_{i,j}(H)$. Let $K$ be a unique non-empty compact subset of $[0,1]$ such that $K=S(K)$. Then, $K$ is a generalized Cantor set, and in terms of the base-7 expansion, we obtain that
@@
and $f_{i,j}([0,1])$ are disjoint. Hence, $K_{1}\subset[0,3/7]$ and $K_{2}\subset[4/7,1]$. In particular, $K_{1}$ and $K_{2}$ are disjoint. Since each $f_{i,j}$ is injective, $f_{i,j}(K_{j})\subsetneq f_{i,j}(K_{1}\cup K_{2})\subset f_{i,j}([0,1])$. Therefore,
@@
In Example 4.5, each $f_{i,j}$ is an affine function. Now we give an example such that all $f_{i,j}$ are non-affine functions.
Example 4.6.
Let $X=[0,1]$ and $d(x,y):=|x-y|$. We consider a graph-directed set on $X$. Let $q=2$. Assume that $|E_{i,j}|=1$ for every $i,j\in\{1,2\}$. We denote $f_{e}$ for $e\in E_{i,j}$ by $f_{i,j}$. Let
$f_{1,1}(x)=\frac{x}{6x+1},\,f_{1,2}(x)=\frac{-x^{2}+2x+4}{7},\,f_{2,1}(x)=\frac{x^{2}+2}{7},\,f_{2,2}(x)=\frac{6-5x}{7-6x}.$
Then, each $f_{i,j}$ is a weak contraction in the sense of Matkowski-Rus. Let $g_{1,1}:=f_{1,1}$, $g_{2,2}:=f_{2,2}$, $g_{1,2}:=f_{1,2}\circ r$, and $g_{2,1}:=f_{2,1}\circ r$. Then, $r\circ g_{1,1}\circ r=g_{2,2}$, $r\circ g_{1,2}\circ r=g_{2,1}$.
We define finite sets $A_{n}$ and $B_{n}$, $n\geq 1$, in the same manner as in Example 4.5. We see that $T^{n}({\bf H}_{0})=(A_{n},B_{n})$ for $n\geq 0$. Let $K_{i},i=1,2$, be non-empty compact subsets of $X=[0,1]$ such that $K_{i}=g_{i,1}(K_{i})\cup g_{i,2}(K_{i}),i=1,2$. Let ${\bf K}:=(K_{1},K_{2})$. Then, we can show that (4.4).
Let $S(H):=\cup_{i,j\in\{1,2\}}f_{i,j}(H)$. Let $K$ be a unique non-empty compact subset of $[0,1]$ such that $K=S(K)$. Since $K_{1}\subset[0,1/7]\cup[4/7,5/7]$ and $K_{2}\subset[2/7,3/7]\cup[6/7,1]$, $K_{1}$ and $K_{2}$ are disjoint. We can show that $K_{1}\cup K_{2}\subsetneq K$.
We remark that $K_{1}=K_{2}=K$ can happen.
Example 4.7.
We keep the same notation as in Example 4.5. Let
$f_{1,1}(x)=\frac{x}{x+1},\,f_{1,2}(x)=\frac{x+1}{2},\,f_{2,1}(x)=\frac{x}{2},\,f_{2,2}(x)=\frac{1}{2-x}.$
Then, each $f_{i,j}$ is a weak contraction in the sense of Matkowski-Rus. We see that
$f_{1,1}([0,1])=f_{2,1}([0,1])=\left[0,\frac{1}{2}\right],\ f_{1,2}([0,1])=f_{2,2}([0,1])=\left[\frac{1}{2},1\right].$
Hence, $T({\bf K})={\bf K}$, where we let $K_{1}=K_{2}=[0,1]$ and ${\bf K}:=(K_{1},K_{2})$. Furthermore, $S([0,1])=[0,1]$.
We can give many examples. [21, Example 3] gives an example different from the above.
@@ 5. Graph-directed de Rham functional equation
In this section, we generalize the framework of the de Rham functional equation to the graph-directed setting. Functional equations indexed by a directed network were recently considered in [22]. Here we deal with the functional equation on $[0,1]$ indexed by a graph-directed finite graph in terms of the graph-directed invariant set and conjugate equations [7, 23]. For simplicity we focus on the case corresponding to $q=2$.
@@
We state several properties of the compatible family $h$. Let
@@
Since $\left|I^{h}_{i}(i_{1},\dots,i_{n})\right|\leq\phi^{n}(1)$ and $\phi$ is a comparison function,
@@
Let $\{f_{i,j}\}_{i,j}$ and $\{g_{i,j}\}_{i,j}$ be two compatible families on $[0,1]$ with a common comparison function $\phi$ under the Euclidean distance [^3] We can replace the Eucliean metric with a different metric by the semi-metric transformer in the above section, as long as $\{f_{i,j}\}_{i,j}$ and $\{g_{i,j}\}_{i,j}$ are contractions in the sense of Matkowski-Rus under the new metric.. For each $i=1,2$, we define functions $\varphi_{i}$, on $D^{f}_{i}$ such that for $i_{1},\dots,i_{n}\in\{1,2\},x_{n}\in\{0,1\}$,
@@
In the same manner, we see that $(i_{k},j_{k})=(2,1)$ for $k\in\{\ell+1,\dots,n\}$ by induction on $k$, and $(x_{n},x_{n}^{\prime})=(1,0)$. By the compatibility condition (5.1) for $\{g_{i,j}\}$, it holds that
@@
On $[0,1]$, we put the topology induced by the Euclidean metric.
@@
(iii) A pair of maps $(\varphi_{1},\varphi_{2})$ satisfies the following system of functional equations:
@@
Assume that $x\in D^{h}_{i}\cap(0,1)$. Then, there exists $n\geq 1$ such that $x\in D^{h}_{i,n}\setminus D^{h}_{i,n-1}$, where we let $D^{h}_{i,0}:=\emptyset$. Then, there exist $i_{1},\dots,i_{n-1}\in\{1,2\}$ such that
@@
Second, we state a connection between the functional equation and the graph-directed set. Put the product topology on $[0,1]^{2}$. Let
@@
Let ${\bf D}_{n}:=(D_{1,n},D_{2,n})$. By induction on $n$, we see that $T^{n}\left({\bf H}_{0}\right)={\bf D}_{n-1}$ for $n\geq 2$. By the arguments in Section 3,
@@
Discrete approximation is a natural approach for analysis for (5.6) and the framework of Theorem 1.1 aligns with the functional equation well.
@@
Acknowledgements  The author would like to give his thanks to a referee for his or her comments. He is grateful to Omer Cantor for comments regarding Remark 2.10. He is also grateful to a production editor for improving the English of the manuscript.